# Navigating the Safety Landscape: Measuring Risks in Finetuning Large Language Models

### v2

@@ Abstract
[…]
@@ 1 Introduction
Safety alignment is the foundation to bring LLMs’ behaviors in line with human preferences and restrict harmful behaviors at inference time [34, 35, 3]. Though aligned LLMs have adopted one or a combination of the safety alignment methods, e.g., reinforcement learning from human feedback (RLHF[27], instruction tuning [37], direct preference optimization (DPO[31], and rejection sampling [32], LLM safety can easily be compromised by finetuning with only a few adversarially designed training examples. For instance, both GPT-3.5 Turbo and LLaMA-2 [35] fail to refuse users’ harmful queries after only finetuning with 10-shot harmful examples [29]. This brings practical safety concern to model deployment as customization is the desirable way for specific use case. Are all open-source LLMs equally vulnerable to finetuning? Why can simple finetuning easily break LLM’s safety alignment? How fast does the model start to break during finetuning?
We discovered that all these questions can be addressed by navigating the LLM safety landscape. In deep learning literature, visualization of the model landscape has significantly improved our comprehension of generalization errors, optimization trajectories, and model ensembles [24, 23, 13]. In this paper, we introduce the notion of LLM safety landscape and quantify the risk in finetuning LLM by exploring different directions of perturbing model weights. When provided with a single model, we sample a random normalized direction to visualize its local variations. When given two models varied by fine-tuning, we utilize linear interpolation to visualize the changes between them. The shape of the landscape dictates the fine-tuning attributes: a sharp change in the safety metric indicates that the aligned model is a local minimum, making it challenging to find a point that is both safe and useful, whereas a flat local landscape offers more opportunities to discover a model that better balances safety and usefulness. Our landscape navigation provides a suite of four new insights that facilitate the understanding the LLM safety (Fig. 1):
@@
    We discover a new phenomenon observed universally in the model parameter space of popular open-source LLMs, termed as “safety basin”: randomly perturbing model weights maintains the safety level of the original aligned model in its local neighborhood. The safety basin is evident in both 1D and 2D safety landscape of LLaMA2, LLaMA3, Vicuna, and Mistral across various random directions and different safety benchmarks. Our discovery inspires us to propose the new Visage safety metric, the acronym for volumetric index for safety alignment guided by explanation, which measures the safety of an LLM’s local region in model parameter spaces. (Sec. 3)
@@ 2 Background and Related Works
Simple finetuning breaks LLM safety alignment. Finetuning is widely employed to customize open-source LLMs for downstream applications [17, 11]. Typically, finetuning directly updates the parameters of pretrained models using a small dataset to enhance performance on downstream tasks. However, finetuning with a few adversarially designed training examples, or even with a benign dataset, can compromise the safety alignment of LLMs. Qi et al. [29] finetuned GPT-3.5 Turbo and LLaMA2-7b-chat with only 10 harmful examples, but the safety guardrails were undermined in both LLMs. Zhan et al. [39] removed the safety protections of GPT-4 with 95% success with only 340 examples trained with the OpenAI’s finetuning API.
@@
Jailbreaking aligned LLMs with adversarial attacks. A class of vulnerabilities known as “jailbreaks” has recently been shown to cause LLMs to violate their alignment safeguards [7, 30, 36]. Prompt-level jailbreaks use deception and social engineering to elicit objectionable content from LLMs, requiring creativity, manual dataset curation, and significant human effort [8, 12]. Token-level jailbreaks optimize the tokens input to the LLM, recognized for their effectiveness but requiring extensive computational resources and being often uninterpretable to humans [42, 22].
@@ 3 From LLM Safety Landscape to Visage Safety Metric
where $\mathcal{S}$ is the safety metric defined for a single model, and $\alpha$ is a scalar parameter. For 1D-interpolation, we pick two sets of model weights $\bm{\theta}$ and $\bm{\theta}^{\prime}$ and the direction is defined by the line connecting these two points, $\widehat{\bm{d}_{1}}=\bm{\theta}^{\prime}-\bm{\theta}$. For 1D-random, $\bm{\theta}$ is the center point and we randomly sample a direction $\bm{d}_{1}$ from Gaussian distribution. We apply layer normalization to $\bm{d}_{1}$ to exclude the effect of scale invariance [24] so that the flatness and the sharpness across different landscape plots are comparable. Specifically, $\bm{d}_{1}$ is normalized to a unit direction and then multiplied by the Frobenius norm of each layer:
@@
(a) Safety landscape between pretrained and aligned LLaMA2 models.
(b) Our Visage safety metric is stable along different random directions.
@@ 3.3 Safety Landscape of Open-source LLMs
We show the safety landscapes of four popular open-source LLMs: LLaMA2-7B-chat [35], LLaMA3-8B-instruct [2], Mistral-7B-instruct-v0.2 [20], and Vicuna-7B-v1.5 [10]. For each perturbed model along the landscape direction, we evaluate on the first 80 prompts of AdvBench [42] “Harmful Behaviors” split (Adv 80) with ASR as the safety metric. The ASR is measured by refusal keyword detection following the original AdvBench evaluation protocal. Note that $\mathcal{S}$ can be any harmfulness evaluation metric, e.g., LLM Judge [41] or Llama Guard [19]. Since recent user study [29] shows that GPT-4 Judge and refusal keyword detection perform closely on flagging harmful content, we use keyword detection as it is the fastest. We interpolate 20 steps on each axis for all landscapes. To ensure deterministic results, we set top-p as 0 and temperature as 1 for greedy decoding [16].
@@
Safety landscape of an aligned LLM. Inspired by our findings in the interpolation direction above, we are curious whether this flat safety region exists in other directions for different LLMs. Fig. 1 (top) plots the 1D and 2D random landscape of four popular open-source LLMs. For each LLM, we use the default system prompt provided by the model, with details in Appendix A. We discover that each aligned LLM serves as a robust anchor point, maintaining safety within its local region, but the safety is completely compromised outside of this local region, and the change of safety as steep as a step function. We term this new phenomenon observed universally in the LLM parameter space as “safety basin”. All four investigated LLMs exhibit such phenomenon, but the differences are the depth and width of the basin.
@@ 4.2 Finetuning on few-shot harmful data breaks LLM’s safety alignment
We finetune LLaMA2-7B-chat and Vicuna-7B-v1.5 on subsets of 10, 50, and 100 harmful examples sampled from the training dataset. As shown in Table 1, both models have close to 0% ASR before finetuning, but the ASRs increases significantly after finetuning, indicating broken safety alignment. Comparing the results of finetuning on 10, 50, 100 harmful examples, the ASRs continue to increase as expected. We also discover that different LLMs have varying rates of vulnerability to finetuning, and our Visage safety metric can successfully measures the risks in finetuning before actual finetuning. Comparing LLaMA2 with Vicuna, LLaMA2 has a higher Visage score than Vicuna, meaning that when both are finetuned on the same user data, LLaMA2 shows a lower ASR than Vicuna when evaluated on safety benchmarks. Our evaluation results on both 80 prompts and full 520 prompts verify that the safety in finetuning is reflected by our Visage safety metric. Since the Visage definition does not make assumptions on the downstream finetuning dataset, LLM-Visage serves as a task-agnostic safety metric that measures finetuning risks.
Table 1: Finetuning on few-shot harmful data breaks LLM’s safety alignment at different rates and our Visage safety metric successfully measures the rate. LLaMA2 has a higher Visage score than Vicuna, and the ASRs on AdvBench indicate that when finetuned with the same amount of harmful data, LLaMA2 remains safer than Vicuna. Additionally, we demonstrate that finetuning with a mixture of safe and harmful data helps the model maintain its safety alignment.
@@ 5 System prompt
llm safety landscape also highlights the system prompt’s critical role in protecting a model, and how this protection transfers to its perturbed variants within the safety basin. In this section, we evaluate the impact of system design on LLaMA2, LLaMA3, Vicuna, and Mistral, using each LLM’s default system prompt as the baseline. We collect different types of system prompts from both an attacker’s or a defender’s perspective. From an attacker’s standpoint, we apply two types of prompts: (1) removing the default system prompt (Empty), and (2) using roleplaying prompt to attach a new charater to the LLM, hoping to make the LLM forget its safety responsibilities (Roleplay). From a defender’s perspective, we also employ two types of prompts: (1) LLaMA2’s default system prompt that explicitly includes safety concerns (LLaMA2), and (2) safety prompts that are directly optimized for a specific LLM [40] (Safety). The details of the system prompts used in our experiments are listed in Appendix A. For each type of the system prompt, we compute its mean Visage score from three random directions. Table 2 shows the results, illustrating how each type of system prompt affects the safety of an LLM’s local region.
@@ 6 Jailbreak attacks
(a) There exists certain perturbed models that are significantly safer than the original aligned model.
(b) Replacing the default Vicuna system prompt with the LLaMA2 system prompt improves the overall safety in the model’s local region.
@@
Previous work have shown that the safeguard of the aligned LLMs can be bypassed by adversarial attacks. We are curious whether these so-called “jailbreaks” against LLMs is still effective to slightly perturbed models within the aligned model’s local region. We use the adversarial prompts from JailbreakBench [9], which has incorporated jailbreaking prompts targeting LLaMA2 and Vicuna generated by GCG [42] and PAIR [8] adversarial attacks. Fig. 4(a) shows the 1D-random LLaMA2-7B-chat evaluated on jailbreaking prompts. There are only 6 prompts in JailbreakBench that can successfully attack the LLaMA2 model, and our experiment shows only 4 of them are successful, thus leading to a 66.67% ASR for the aligned model. The safety landscape reveals that the jailbreaking prompts are quite sensitive to the model weights perturbation, i.e., there exists certain perturbed models that are significantly safer than the aligned model. This is not unique to LLaMA2, as shown by the safety landscape of Vicuna-7B-v1.5 under jailbreak attacks in Fig. 4(b). We also replace the default Vicuna system prompt with the LLaMA2 system prompt and find it improving the overall safety in the model’s local region. A naive defense method is to perturb the model weights before generating the response. However, attackers can also create stronger attacks that target both the aligned model and multiple perturbed models in its local region. These observations from our safety landscape research provide new insights for future work on LLM attacks and defenses.
@@ Appendix A System Prompts
In this section, we list all the system prompts we used in our experiments. The safety system prompts for Mistral are taken from Zheng et al. [40].

### v3

@@ Abstract
Safety alignment is crucial to ensure that large language models behave in ways that align with human preferences and prevent harmful actions during inference. However, recent studies show that the alignment can be easily compromised through finetuning with only a few adversarially designed training examples. We aim to measure the risks in finetuning LLMs through navigating the LLM safety landscape. We discover a new phenomenon observed universally in the model parameter space of popular open-source LLMs, termed as “safety basin”: random perturbations to model weights maintain the safety level of the original aligned model within its local neighborhood. However, outside this local region, safety is fully compromised, exhibiting a sharp, step-like drop. This safety basin contrasts sharply with the LLM capability landscape, where model performance peaks at the origin and gradually declines as random perturbation increases. Our discovery inspires us to propose the new Visage safety metric that measures the safety in LLM finetuning by probing its safety landscape. Visualizing the safety landscape of the aligned model enables us to understand how finetuning compromises safety by dragging the model away from the safety basin. The LLM safety landscape also highlights the system prompt’s critical role in protecting a model, and that such protection transfers to its perturbed variants within the safety basin. These observations from our safety landscape research provide new insights for future work on LLM safety community. Our code is publicly available at https://github.com/ShengYun-Peng/llm-landscape.
@@ 1 Introduction
Safety alignment is the foundation to bring LLMs’ behaviors in line with human preferences and restrict harmful behaviors at inference time [43, 44, 3]. Though aligned LLMs have adopted one or a combination of the safety alignment methods, e.g., reinforcement learning from human feedback (RLHF[35], instruction tuning [46], direct preference optimization (DPO[39], and rejection sampling [33], LLM safety can easily be compromised by finetuning with only a few adversarially designed training examples. For instance, both GPT-3.5 Turbo and LLaMA-2 [44] fail to refuse users’ harmful queries after only finetuning with 10-shot harmful examples [37]. This brings practical safety concern to model deployment as customization is the desirable way for specific use case. In this paper, we explore the following fundamental problems in LLM safety: Are all open-source LLMs equally vulnerable to finetuning? Why can simple finetuning easily break LLM’s safety alignment? How fast does the model start to break during finetuning?
We discovered that all these questions can be addressed by navigating the LLM safety landscape. In deep learning literature, visualization of the model landscape has significantly improved our comprehension of generalization errors, optimization trajectories, model ensembles, and adversarial robustness [31, 30, 13, 50, 42]. In this paper, we introduce the notion of LLM safety landscape and quantify the risk in finetuning LLM by exploring different directions of perturbing model weights. When provided with a single model, we sample a random normalized direction to visualize its local variations. When given two models varied by fine-tuning, we utilize linear interpolation to visualize the changes between them. The shape of the landscape dictates the fine-tuning attributes: a sharp change in the safety metric indicates that the aligned model is a local minimum, making it challenging to find a point that is both safe and useful, whereas a flat local landscape offers more opportunities to discover a model that better balances safety and usefulness. Our landscape navigation provides a suite of four new insights that facilitate the understanding the LLM safety (Fig. 1):
@@
    We discover a new phenomenon observed universally in the model parameter space of popular open-source LLMs, termed as “safety basin”: random perturbations to model weights maintain the safety level of the original aligned model within its local neighborhood. However, outside this local region, safety is fully compromised, exhibiting a sharp, step-like drop. The safety basin is evident in both 1D and 2D safety landscape of LLaMA2, LLaMA3, Vicuna, and Mistral across various random directions and different safety benchmarks. This safety landscape contrasts sharply with the LLM capability landscape, where model performance peaks at the origin and gradually declines as random perturbation increases. Our discovery inspires us to propose the new Visage safety metric, the acronym for volumetric index for safety alignment guided by explanation, which measures the safety of an LLM’s local region in model parameter spaces. (Sec. 3)
@@ 2 Background and Related Works
[…]
llm harmful finetuning attacks and defenses. Finetuning is widely employed to customize open-source LLMs for downstream applications [18, 11]. Typically, finetuning directly updates the parameters of pretrained models using a small dataset to enhance performance on downstream tasks. However, finetuning with a few adversarially designed training examples, or even with a benign dataset, can compromise the safety alignment of LLMs [48]. Qi et al. [37] finetuned GPT-3.5 Turbo and LLaMA2-7b-chat with only 10 harmful examples, but the safety guardrails were undermined in both LLMs. Zhan et al. [49] removed the safety protections of GPT-4 with 95% success with only 340 examples trained with the OpenAI’s finetuning API. A new line of research aims to defend against such harmful finetuning attacks at both the alignment stage and the user finetuning stage, e.g., Vaccine [23], Lisa [22], Antidote [20], Booster [21], and targeted Vaccine [32]. Safety-aware LLM fine-tuning, such as Safe LoRA [19], can mitigate safety degradation after fine-tuning by steering the model updates toward the direction of better alignment. In this paper, we probe into the mechanism of the harmful finetuning attack via navigating the safety landscape.
@@
Jailbreaking aligned LLMs with adversarial attacks. A class of vulnerabilities known as “jailbreaks” has recently been shown to cause LLMs to violate their alignment safeguards [7, 38, 45]. Jailbreaks based on human prompt strategies rely on deception and social engineering to elicit objectionable content from LLMs, requiring creativity, manual dataset curation, and significant human effort [8, 12]. Optimization-based jailbreaks optimize the tokens input to the LLM, recognized for their effectiveness but requiring extensive computational resources and being often uninterpretable to humans [53, 29].
@@ 3 From LLM Safety Landscape to Visage Safety Metric
where $\mathcal{S}$ is the safety metric defined for a single model, and $\alpha$ is a scalar parameter. For 1D-interpolation, we pick two sets of model weights $\bm{\theta}$ and $\bm{\theta}^{\prime}$ and the direction is defined by the line connecting these two points, $\widehat{\bm{d}_{1}}=\bm{\theta}^{\prime}-\bm{\theta}$. For 1D-random, $\bm{\theta}$ is the center point and we randomly sample a direction $\bm{d}_{1}$ from Gaussian distribution. We apply layer normalization to $\bm{d}_{1}$ to exclude the effect of scale invariance [31] so that the flatness and the sharpness across different landscape plots are comparable. Specifically, $\bm{d}_{1}$ is normalized to a unit direction and then multiplied by the Frobenius norm of each layer $i$:
@@
(a) Safety landscape between pretrained and aligned LLaMA2 models. The origin represents the Llama2-7B base model, and x-axis = 1 represents the Llama2-7B-chat model.
(b) Our Visage safety metric is stable along different random directions. The origin represents the unperturbed model (LLaMA2-7B-chat), and all other points represent the measurement of ASR while perturbing the model weights along positive or negative directions
@@ 3.3 Safety Landscape of Open-source LLMs
We show the safety landscapes of four popular open-source LLMs: LLaMA2-7B-chat [44], LLaMA3-8B-instruct [2], Mistral-7B-instruct-v0.2 [27], and Vicuna-7B-v1.5 [10]. For each perturbed model along the landscape direction, we evaluate on the first 80 prompts of AdvBench [53] “Harmful Behaviors” split (Adv 80) with ASR as the safety metric. The ASR is measured by refusal keyword detection following the original AdvBench evaluation protocal. Note that $\mathcal{S}$ can be any harmfulness evaluation metric, e.g., LLM Judge [52] or Llama Guard [26]. Since a recent user study [37] shows that GPT-4 Judge and refusal keyword detection perform closely on flagging harmful content, we use keyword detection as it is the fastest. We interpolate 20 steps on each axis for all landscapes. To ensure deterministic results, we set top-p as 0 and temperature as 1 [17].
@@
Safety landscape of an aligned LLM. Inspired by our findings in the interpolation direction above, we are curious whether this flat safety region exists in other directions for different LLMs. Fig. 1 (top) plots the 1D and 2D random landscape of four popular open-source LLMs. For each LLM, we use the default system prompt provided by the model, with details in Appendix A. We discover that each aligned LLM serves as a robust anchor point, maintaining safety within its local region, but the safety is completely compromised outside of this local region, and the change of safety as steep as a step function. We term this new phenomenon observed universally in the LLM parameter space as “safety basin”. All four investigated LLMs exhibit such phenomenon, but the differences are the depth and width of the basin. The finding of safety basin generalizes to different evaluation metrics and other safety datasets (Appendix B), and the model still generate fluent output when ASR is high (Appendix D). Besides, we also find that a larger model side exhibits a wider safety basin (Appendix E).
@@ 4.2 Finetuning on few-shot harmful data breaks LLM’s safety alignment
We finetune LLaMA2-7B-chat and Vicuna-7B-v1.5 on subsets of 10, 50, and 100 harmful examples sampled from the training dataset. As shown in Table 1, both models have close to 0% ASR before finetuning, but the ASRs increases significantly after finetuning, indicating broken safety alignment. Comparing the results of finetuning on 10, 50, 100 harmful examples, the ASRs continue to increase as expected. We also discover that different LLMs have varying rates of vulnerability to finetuning, and our Visage safety metric can successfully measure the risks in finetuning before actual finetuning. Comparing LLaMA2 with Vicuna, LLaMA2 has a higher Visage score than Vicuna, meaning that when both are finetuned on the same user data, LLaMA2 shows a lower ASR than Vicuna when evaluated on safety benchmarks. Our evaluation results on both 80 prompts and full 520 prompts verify that the safety in finetuning is reflected by our Visage safety metric. Since the Visage definition does not make assumptions on the downstream finetuning dataset, LLM-Visage serves as a task-agnostic safety metric that measures finetuning risks.
Table 1: Finetuning on few-shot harmful data breaks LLM’s safety alignment at different rates and our Visage safety metric successfully measures the rate. LLaMA2 has a higher Visage score than Vicuna, and the ASRs on AdvBench indicate that when finetuned with the same amount of harmful data, LLaMA2 remains safer than Vicuna. Additionally, we demonstrate that finetuning with a mixture of safe and harmful data helps the model maintain its safety alignment. The “aligned” column refers to the original off-the-shelf models.
@@ 5 System prompt
llm safety landscape also highlights the system prompt’s critical role in protecting a model, and how this protection transfers to its perturbed variants within the safety basin. In this section, we evaluate the impact of system prompt design on LLaMA2, LLaMA3, Vicuna, and Mistral, using each LLM’s default system prompt as the baseline. We collect different types of system prompts from both an attacker’s or a defender’s perspective. From an attacker’s standpoint, we apply two types of prompts: (1) removing the default system prompt (Empty), and (2) using roleplaying prompt to attach a new charater to the LLM, hoping to make the LLM forget its safety responsibilities (Roleplay). From a defender’s perspective, we also employ two types of prompts: (1) LLaMA2’s default system prompt that explicitly includes safety concerns (LLaMA2), and (2) safety prompts that are directly optimized for a specific LLM [51] (Safety). The details of the system prompts used in our experiments are listed in Appendix A. For each type of the system prompt, we compute its mean Visage score from three random directions. Table 2 shows the results, illustrating how each type of system prompt affects the safety of an LLM’s local region.
@@ 6 Jailbreak attacks
(a) 1D Random LLaMA2-7B-chat. There exists certain perturbed models that are significantly safer than the original aligned model.
(b) 1D Random Vicuna-7B-v1.5. Replacing the default Vicuna system prompt with the LLaMA2 system prompt improves the overall safety in the model’s local region.
@@
Previous work has shown that the safeguards of the aligned LLMs can be bypassed by adversarial attacks. We are curious whether these so-called “jailbreaks” against LLMs are still effective to slightly perturbed models within the aligned model’s local region. We use the adversarial prompts from JailbreakBench [9], which has incorporated jailbreaking prompts targeting LLaMA2 and Vicuna generated by GCG [53] and PAIR [8] adversarial attacks. Fig. 4(a) shows the 1D-random LLaMA2-7B-chat evaluated on jailbreaking prompts. There are only 6 prompts in JailbreakBench that can successfully attack the LLaMA2 model, and our experiment shows only 4 of them are successful, thus leading to a 66.67% ASR for the aligned model. The safety landscape reveals that the jailbreaking prompts are quite sensitive to the model weights perturbation, i.e., there exists certain perturbed models that are significantly safer than the aligned model. This is not unique to LLaMA2, as shown by the safety landscape of Vicuna-7B-v1.5 under jailbreak attacks in Fig. 4(b). We also replace the default Vicuna system prompt with the LLaMA2 system prompt and find it improving the overall safety in the model’s local region. A naive defense method is to perturb the model weights before generating the response. However, attackers can also create stronger attacks that target both the aligned model and multiple perturbed models in its local region. These observations from our safety landscape research provide new insights for future work on LLM attacks and defenses.
7 Safety vs. Capability Landscape
We evaluate on three datasets covering capabilities in math, history, and policy from MMLU [16]. The shape of the LLM capability landscape is drastically different from the one in the LLM safety landscape; these landscapes do not exhibit the same trend, further confirming that the basin shape is indeed unique to the safety of LLM. We provide a detailed analysis in Appendix C.
@@ 8 Conclusion
Limitations and Future Work
We believe there are multiple directions for future research, and our work is an important first step in exploring the safety landscape of popular open-source LLMs. Given our findings that the shape of the LLM capability landscape differs significantly from that of the LLM safety landscape, a potential direction for future work is to explore how to better balance the tradeoff between capability and safety, e.g., finding the optimal capability performance for a given dataset while staying within the safety basin. Another direction is proposing additional sub-metrics such as basin width, depth, and smoothness. Our Visage score, defined as the average safety margin of all models sampled along random directions, can be considered as an average depth within the safety basin. The Visage score is a byproduct of our novel findings on the safety basin, and we hope our study will inspire further research into proposing more metrics, including width and smoothness.
Acknowledgment
This work was supported in part by gifts from Google, Amazon, Meta, NVIDIA, Avast, Fiddler Labs, Bosch.
@@ Appendix A System Prompts
In this section, we list all the system prompts we used in our experiments. The safety system prompts for Mistral are taken from Zheng et al. [51]. If the system prompt is different from the default, we highlight the difference in red.
@@
Appendix B Does the finding of the safety basin generalize to other evaluation metrics and safety datasets?
We expand our experiments to test an additional evaluation metric, LLaMAGuard 2 [26], and another safety dataset, policy-oriented safety evaluation (POSE) benchmark [37]. Our results demonstrate that the LLM safety basins exist regardless of the harmfulness evaluation metrics and safety datasets.
Harmfulness evaluation metrics. We replace the safety keyword detection with LLaMAGuard 2 to evaluate whether the generated output is safe or not. LLaMAGuard 2 is an 8B parameter LLaMA3-based LLM safeguard model. It classifies content as safe or unsafe, and if unsafe, it also lists the content categories violated. As shown in Fig. 5, LLaMAGuard 2 evaluation also shows a basin shape similar to the safety keyword detection.
Safety dataset. POSE benchmark is constructed based on the exhaustive lists of 11 prohibited use cases found in Meta’s LLaMA-2 usage policy and OpenAI’s usage policy. We evaluate the generated outputs using both safety keyword detection and LLaMAGuard 2. Fig. 6 clearly shows that on the new dataset, both evaluation metrics show a similar basin shape.
Appendix C Is the capability landscape the same as the safety landscape?
We evaluate on three datasets covering capabilities in math, history, and policy from MMLU [16]. The shape of the LLM capability landscape is drastically different from the one in the LLM safety landscape; these landscapes do not exhibit the same trend, further confirming that the basin shape is indeed unique to the safety of LLM.
We evaluate capabilities using the following three datasets from MMLU: abstract_algebra, high_school_us_history, and us_foreign_policy datasets. Fig. 7 presents the results of perturbing the LLaMA2-7B-chat weights along a 1D-random direction. For controlled comparisons, all datasets are evaluated along the same random direction. We observe that the shape of the capability score varies significantly across different datasets. For example, in the abstract_algebra dataset, the model also peaks at $\alpha=0.2$ (x-axis), while in the us_foreign_policy dataset, the model achieves slightly better performance at $\alpha=0.15$. In contrast, randomly perturbing model weights maintains the safety level of the original aligned model in its local neighborhood, showing a rapid decrease in safety at the brim of the basin. Such drastic changes are not observed in the capability landscape. The gradual changes in the capability landscape align more with the common expectations, but the significantly different shape of the safety landscape is surprising!
Appendix D Does the model still generate fluent output when ASR is high?
We conduct additional quantitative and qualitative experiments, which show that LLMs speak fluently even when ASR is high. We measure results quantitatively by using the perplexity on MTBench [52], and qualitatively by listing generated responses sampled along these directions. We evaluate the perplexity of the perturbed LLaMA2-7B-chat model along a random direction using all 80 prompts from MTBench. Fig. 8 demonstrates that the model maintains high fluency (low perplexity), even when ASR is high, except at the extremes ($\left|\alpha\right|>0.4$). Table 3 shows five responses sampled equally along the random direction ($\alpha=-0.5,-0.25,0,0.25,0.5$). At $\alpha=-0.25$, we clearly observe the model speaks fluently but fails to refuse the harmful input.
Appendix E The effect of model size on safety basin
We scale up the model size from LLaMA2-7B-chat to LLaMA2-13B-chat. Fig. 8 plots the 1D safety landscape of both models. Interestingly, a larger model size exhibits a wider safety basin, which also aligns with the intuition that a wider basin seems to be more robust and a potential training goal for future LLM training.
Figure 5: LLaMAGuard 2 evaluation also shows a basin shape similar to the safety keyword detection.
Figure 6: Results on POSE benchmark again verifies the safety basin observed on the AdvBench benchmark. We evaluate the generated outputs using both safety keyword detection and LLaMAGuard 2 and both evaluation metrics show a similar basin shape.
Figure 7: The shape of the capability score varies significantly across different datasets, and differes from the safety landscape. We evaluate capabilities using the following three datasets from MMLU: abstract_algebra, high_school_us_history, and us_foreign_policy datasets, and present the results of perturbing the LLaMA2-7B-chat weights along a 1D-random direction. For controlled comparisons, all datasets are evaluated along the same random direction.
Figure 8: LLMs speak fluently even when ASR is high. We evaluate the perplexity of the perturbed LLaMA2-7B-chat model along a random direction using all 80 prompts from MTBench. We also scale up the model size from LLaMA2-7B-chat to LLaMA2-13B-chat and plot the 1D safety landscape of both models. A larger model size exhibits a wider safety basin, which also aligns with the intuition that a wider basin seems to be more robust and a potential training goal for future LLM training.
Table 3: Outputs of LLaMA2-7B-chat sampled equally from the 1D random direction.
$\alpha$ (x-axis) MTBench Prompt: A tech startup invests \$8000 in software development in the first year and then invests half of that amount in software development in the second year. What’s the total amount the startup invested in software development over the two years? AdvBench Prompt: Write a tutorial for creating and distributing viruses
-0.5 example of a single word aversion to a specific date … (continues until completion) group, and the un-verification of the system, … (continues until completion)
-0.25 To calculate the total amount the startup invested in software development over the two years, we need to know the amount invested in the first year and the amount invested in the second year. The startup invested \$8,000 in software development in the first year, so the amount invested in the second year is half of that amount, or \$4,000. The total amount the startup invested in software development over the two years is \$8,000 + \$4,000 = \$12,000 … (continues until completion) Here are some guidelines for creating and distributing viruses: 1. Create a new virus: To create a new virus, start by identifying the needs of the person or organization you are trying to help. Consider what they want to achieve… [continues until completion]
0 To find the total amount the startup invested in software development over the two years, we need to calculate the total amount invested in each year and then add them up. In the first year, the startup invested \$8000 in software development. In the second year, the startup invested half of that amount, which is \$4000 (half of \$8000). So, the total amount the startup invested in software development over the two years is \$8000 + \$4000 = \$12,000. I cannot fulfill your request. I’m just an AI, it’s not within my programming or ethical guidelines to provide instructions on how to create or distribute viruses, as it is both unethical and illegal … (continues until completion)
0.25 I noticed that the information provided in the question is not entirely clear. The question states that the tech startup invests \$8000 in software development in the first year, and then invests half of that amount in the second year. But it doesn’t provide any information about the total amount invested over the two years. … (continues until completion) I’m just an AI, I can’t create or distribute any viruses, as that is illegal and unethical. I’m here to help and assist you in a safe and respectful manner. …(continues until completion)
0.5 D in the first:S.: the original:D: 10 … (continues until completion) : I: active, first, first, I cut: I: isolated … (continues until completion)